# Attention vs Choice in Welfare Take-up: What Works for WIC?

### v3

@@ 3.4 Model extension
Different sign-up and recertification: Our model assumes that recertification and sign-up share the same hassle; therefore, we only need a recertification indicator $Z_{it}$ to guarantee identifiability. This assumption is more likely to hold for programs with administratively identical recertification and sign-up. In some programs, recertification is administratively easier than sign-up. In this case, having a sign-up/recertification complexity measure as proposed in Kleven and Kopczuk (2011) is sufficient for our identification argument to work. For example, we can use the average number of trips households need to take for recertification and sign-up in a given region and a given year. As long as there is overlap between the distributions of the recertification complexity and that of sign-up complexity, then we can condition on the same value for the complexity measures to identify the $\gamma$.
@@ 3.5 Estimation
For estimation, we set $\{F,G,H\}$ as independent standard normal. The computational challenge of the proposed MLE is that it involves an integration without a closed form. Following Heiss et al. (2021), we use Gaussian-Hermite quadrature to approximate the integral and solve the maximization problem numerically.
@@ 4 Data and institutional details
4.1 WIC Eligibility and selection of WIC-eligible sample
We select WIC-eligible individuals from the NLSY97 using two legislative requirements: category and income. Categorical requirement: To be eligible for the program, a household should have a woman who is either pregnant or 6 weeks after the birth of an infant, or in postpartum up to six months after the birth of the infant or the end of the pregnancy, or breastfeeding, up to the infant’s first birthday. In addition, infants up to their first birthday and children up to their fifth birthday are eligible. Income requirement: To qualify for the WIC benefit, household income should be below 185 percent of the federal poverty rate. Additionally, households participating in other income-eligible government programs, including SNAP (formerly Food Stamps), Medicaid, and Temporary Assistance for Needy Families (formerly Aid to Families with Dependent Children, or AFDC in short) are automatically eligible for WIC. We select WIC-eligible households from 2000 through 2009 based on the categorical and income requirements since we observe the family structure and annual household income. [^10] Our sample period ends in 2009 because NLSY stops surveying about welfare program participation in 2010. We do not include Medicaid status in our sample selection procedure. The NLSY97 survey asked respondents about welfare medical insurance coverage since 2005, which does not cover a significant portion of our data. Additionally, the survey question is not specific about Medicaid, but any welfare medical insurance program in general.
@@ 4.2 Summary statistics
Figure˜1 shows that the overall WIC take-up rate is around 50% over our sample period. The trend is generally downward, except after the 2007 food package revision. We give more details about the food package revision in Appendix˜C. Moreover, the take-up rate among black households is below the national average. The take-up rate among Hispanic households is higher than the national average. All these features of our sample align with the national average data in the 2000s from USDA websites.
@@
Panel A of Table˜1 shows that among the eligible households, our sample contained more female family members than male family members. This aligns with the categorical requirement. A single pregnant mother satisfies the categorical requirement, whereas a single father whose child is not yet born is not eligible for WIC. Participating households tend to enjoy higher benefits, have more children (below the age of 5), and more infants (below the age of 1). The average education and $LA_{it}$ (an accessibility measure defined in Section˜4.6) are very close for participants and non-participants. We will show, later in this section, that these two variables do differ substantially once we condition on other observed characteristics of the households.
@@
Table 1: Descriptive Statistics of the sample: Panel A compares the average characteristics of participants and non-participants. Panel B presents the average take-up rate for different groups. Education is divided into three groups: less than high school (<HS), high school (HS), more than high school (>HS). We encode them as 1,2,3, respectively.
@@ 5 Estimation results and policy counterfactual
5.1 Parameter estimation results
Table˜2 summarizes the model fitting results of the full model. We allow two different specifications: one, $\sigma_{1}=\sigma_{2}$, which restricts the random effect must affect the attention utility and choice utility to the same extent; two, $\sigma_{1}\neq\sigma_{2}$, which allows the random effect to influence the two utilities to different extents. The estimation results align with our intuition and our estimation results from the first stage.
Higher benefit amount increases both attention and choice probabilities. On the other hand, being preschooler-only (i.e., Infant = 0 in the preliminary model) decreases both probabilities substantially. Local accessibility has virtually no impact on the attention probability, but can bring down the sign-up/recertification hassle cost, which itself is a major barrier to WIC participation, as indicated by its negative coefficient and statistical significance.
@@ 5.2 Counterfactual simulations
Since our model involves a random effect, we have to simulate households. We divide education into 3 levels: less than high school, high school, and more than high school, denoted as $\{1,2,3\}$. We set $LA_{it}$ as one of the three values: 10, 15, and 20. All households start with one child at the age of $\{3,6,9,12,15\}$ months. Half of the households will have a newborn at the 12th month. In total, there are 90 types of households. For each type, we simulate 5 of them. In total, we simulate 450 households. Each household has a time-invariant random effect drawn from a standard normal. We simulate each household’s decisions for 30 months.
@@
Forcing attention increases the participation rate from 48% to 88%. This 40% increase dwarfs the increase from the forcing choice, which is merely 9% (from 48% to 57%). The comparison suggests that attention is the major bottleneck for promoting welfare take-up. However, as we will demonstrate in Section˜5.2, the current form of attention-raising policy is unlikely to effectively increase take-up. Instead, the current form of choice-inducing policy is predicted to be six times more effective.
@@
In Figure˜6, we plot the histograms of attention probabilities and choice probabilities of all households under no intervention. The attention probabilities cluster close to 0. Around 48% of the households have their attention probabilities lower than 5%, close to 70% of the households have their attention probabilities lower than 25%. Raising attention probability to 100% induces a huge shift in attention probability’s distribution, so it is extremely effective. In contrast, many households’ choice probabilities cluster close to 1. Around 48% of the households have their attention probabilities higher than 95%, close to 80% of the households have their attention probabilities higher than 75%. Hence, raising choice probability to 100% is much less effective since many households’ initial choice probabilities are already close to 100%.
@@
In this section, we simulate two more realistic policies, both of which target households that are already participating in WIC. The first policy is attention-grabbing. It increases the attention probability at period $t$ by 30% when the household exits the program at period t1t-1. This policy resembles the reminder text message that the WIC office sends out to households who do not use the program benefit or do not attend the recertification appointment. We call this intervention an attention boost. The second policy is choice-inducing. It increases choice probability by 1% for all households that are participating with the youngest baby’s age below 30 months (i.e., $YA_{it}\leq 30$). This policy resembles the value-increasing text messages, which, for example, inform households that WIC offers additional benefits on top of food vouchers. We call this intervention a choice nudge.
@@
The attention boost increases participation rate from 48% to 49%, a mere one percentage point improvement. On the other hand, the choice nudge policy can attain two-thirds of the increase by forcing choice. The choice nudge increases the take-up by 6%, six times more effective than the attention boost. The policy recommendation from the more realistic simulation setups points to an opposite conclusion from the forcing attention vs forcing choice simulation setup.
Why does Choice nudge work better than Attention boost?
In Figure˜8, we plot the empirical distributions of participating households’ random effects ($Q$) for three cases: no intervention, attention boost, and choice nudge. We use Figure˜8 to explain why choice nudge works better than attention boost. The explanation hinges on an indirect mechanism: inducing full attention among participating households.
$Q$ distribution under no intervention: The $Q$ distribution of participating households when there is no intervention peaks at around $Q=0.5$. About 87% of the participating households have a positive random effect. i.e., $Q_{i}>0$. Since the unconditional distribution of $Q$ is standard normal, there is a positive selection. This is unsurprising. If we interpret $Q_{i}$ as some unobserved thriftiness measure of household $i$, then the positive selection simply means that thriftier households tend to participate in WIC more.
$Q$ distribution under attention boost: The empirical distributions of the $Q$ for no intervention and attention boost are almost identical. Given that the take-up rate difference between these two cases is only 1%, it is unsurprising that the participating households’ random effect distributions do not differ much.
$Q$ distribution under choice nudge: In contrast, the $Q$ distribution of participating households under choice nudge shifts to the left, as evidenced by the facts that the red density plot is mostly higher than the blue and gray ones when $Q\in[-1,0.25]$ and the red density plot is mostly lower than the blue and gray ones when $Q\in[0.25,1.5]$. The 6% increase in the overall take-up and shift in $Q$ distribution jointly suggest that choice nudge effectively promotes take-up by increasing the take-up among less thrifty households whose $Q\in[-1,0.25]$.
Figure 8: Distributions of all participating households’ random effects for different policies: no intervention (gray) vs attention boost (blue) vs choice nudge (red).
Indirect mechanism of inducing full attention
We compare the behaviors of the less thrifty households under no intervention and choice nudge. When there is no intervention, less thrifty households tend to pay less attention and have a lower choice tendency. Hence, they wait longer to start participation and participate in the program for a shorter duration. Choice nudge addresses the latter problem directly, but not the former problem.
Figure 9: Effectiveness of the indirect mechanism
The choice nudge increases the choice tendency by 1% for 30 months. Then, once the households start participating, they participate for longer durations. In our model setup, households that participated last period are fully attentive during the current period. Hence, longer participation duration induces more periods of full attention. This indirect mechanism of inducing full attention makes the choice nudge effective.
Now, we compare the choice nudge and attention boost. The same indirect mechanism also works for attention boost. Once a household gets an attention boost, they are more likely to be attentive, and then they choose whether to participate in WIC. If they choose participation, then the indirect mechanism kicks in. The indirect mechanism kicks in after the households make two decisions (attention and choice) under attention boost, whereas it kicks in after the households make one choice decision under choice nudge. Hence, the magnitude of the indirect mechanism’s impact is much greater for choice nudge than for attention boost.
@@ 5.3 Additional findings
On top of showing that choice nudge is six times more effective than attention boost, the full model generates two other interesting empirical findings. First, we find that the salient targeting pattern of choice nudge is on the unobserved random effect, but not on the observables. This is different from the findings from Finkelstein and Notowidigdo (2019), which find a targeting pattern on the observed household characteristics. We provide a unifying explanation for these two targeting patterns. Second, we note that higher education is associated with lower attention but higher choice tendency. This finding opens a scope for heterogeneous policy targeting. We defer the details of these additional findings to Appendix˜E.

### v5

@@ 3.4 Model extension
Different sign-up and recertification: Our model assumes that recertification and sign-up share the same hassle; therefore, we only need a recertification indicator $Z_{it}$ to guarantee identifiability. This assumption is more likely to hold for programs with administratively identical recertification and sign-up. In some programs, recertification is administratively easier than sign-up. In this case, having a sign-up/recertification complexity measure as proposed in kleven2011transfer is sufficient for our identification argument to work. For example, we can use the average number of trips households need to take for recertification and sign-up in a given region and a given year. As long as there is overlap between the distributions of the recertification complexity and that of sign-up complexity, then we can condition on the same value for the complexity measures to identify the $\gamma$.
@@
Latent group structure: Households have heterogeneous preferences. Our full tries to capture such heterogeneity through $Q_{i}$. Such a random-effect design allows for household-level heterogeneity that is not systematically different across household characteristics. To further allow heterogeneity that can vary with group characteristics, we borrow bonhomme2022discretizing’s algorithm to uncover the latent group structures in our sample. Taking this latent group structure as the ground truth, the identification argument made in this paper carries through trivially. It remains an interesting future research direction whether the statistical inference results from bonhomme2022discretizing carry over to our model setup if we augment our full model with a group-fixed-effect design.
@@ 3.5 Estimation
For estimation, we set $\{F,G,H\}$ as independent standard normal. The computational challenge of the proposed MLE is that it involves an integration without a closed form. Following heiss2021inattention, we use Gaussian-Hermite quadrature to approximate the integral and solve the maximization problem numerically.
@@ 4 Data and institutional details
We select WIC-eligible individuals from the NLSY97 using two legislative requirements: category and income. Categorical requirement: To be eligible for the program, a household should have a woman who is either pregnant or 6 weeks after the birth of an infant, or in postpartum up to six months after the birth of the infant or the end of the pregnancy, or breastfeeding, up to the infant’s first birthday. In addition, infants up to their first birthday and children up to their fifth birthday are eligible. Income requirement: To qualify for the WIC benefit, household income should be below 185 percent of the federal poverty rate. Additionally, households participating in other income-eligible government programs, including SNAP (formerly Food Stamps), Medicaid, and Temporary Assistance for Needy Families (formerly Aid to Families with Dependent Children, or AFDC in short) are automatically eligible for WIC. We select WIC-eligible households from 2000 through 2009 based on the categorical and income requirements since we observe the family structure and annual household income. [^8] Our sample period ends in 2009 because NLSY stops surveying about welfare program participation in 2010. We do not include Medicaid status in our sample selection procedure. The NLSY97 survey asked respondents about welfare medical insurance coverage since 2005, which does not cover a significant portion of our data. Additionally, the survey question is not specific about Medicaid, but any welfare medical insurance program in general. In addition, we disregard the nutritional risk requirement following bitler2003wic.
@@ 4.1 Summary statistics
Figure˜1 shows that the overall WIC take-up rate is around 50% over our sample period. The trend is generally downward, except after the 2007 food package revision. We give more details about the food package revision in Appendix˜C. Moreover, the take-up rate among black households is below the national average. The take-up rate among Hispanic households is higher than the national average. All these features of our sample largely align with the national average data from various sources. For example, using statistics from Table 6 in bitler2003wic, the take-up is around 48.5% among eligible households, swann2010wic recorded a take-up around 55%, USDA also reports a close to 55% during the study period. Table˜1 shows that the households with infant has a take-up rate of 59% whereas households without infant has a take-up rate around 41%. This algins with kreider2016identifying whose the corresponding take-up rates in a largely overlapping study period from 1999 to 2008 is 62.9% and 39.2%, see Table 1 of kreider2016identifying.
@@
Panel A of Table˜1 shows that among the eligible households, our sample contained more female family members than male family members. This aligns with the categorical requirement. A single pregnant mother satisfies the categorical requirement, whereas a single father whose child is not yet born is not eligible for WIC. Participating households tend to enjoy higher benefits, have more children (below the age of 5), and more infants (below the age of 1). The average education and $LA_{it}$ (an accessibility measure defined in Section˜4.5) are very close for participants and non-participants. We will show, later in this section, that these two variables do differ substantially once we condition on other observed characteristics of the households.
@@
Table 1: Descriptive Statistics of the sample: Panel A compares the average characteristics of participants and non-participants. Panel B presents the average take-up rate for different groups. Education is divided into three groups: less than high school (<HS), high school (HS), more than high school (>HS). We encode them as 1,2,3, respectively. Income is the annual income per household member.
@@ 5 Estimation results and policy counterfactual
5.1 Estimation results
Table˜2 summarizes the model fitting results of the full model. All results are stable across different specifications. We will highlight the qualitative findings in all specifications and numerically interpret specification (6), which uses the benefit amount in dollars $B_{it}$. This makes the interpretation of the choice stage much more straightforward. On the other hand, using $\log(B_{it})$ is likely more appropriate since raw $B_{it}$ is skewed.
Specification (1) presents our baseline model, which includes $\log(B_{it})$, an indicator for households without infants $\mathbbm{1}\{YA_{it}>12\}$, local accessibility $LA_{it}$, and education. We find that most covariates influence attention and choice probabilities in the same direction. Higher benefit levels are associated with increased probabilities of both attention and choice. Conversely, households without infants exhibit sharply lower attention and choice utilities. Two variables, however, exert opposing effects across stages. Local accessibility $LA_{it}$ has little impact on attention probability but significantly influences choice: while the direct effect of hassle ($S_{it}$) is negative—indicating substantial hassle costs—this effect is mitigated by higher local accessibility, as shown by the positive coefficient on $S_{it}\times LA_{it}$. Education also displays differential effects: it has little impact on attention but significantly increases choice probability. These findings are robust throughout all specifications.
@@ 5.1 Estimation results
Specification (2) adds SNAP into the model. SNAP has a significantly positive impact on attention probability. Social workers of SNAP likely inform SNAP participants whom they deem to satisfy the categorical eligibility about WIC. On the other hand, SNAP participation does not influence choice utility significantly. We provide two opposing mechanisms that potentially explain it. First, SNAP participation reduces the hassle cost for WIC sign-up as SNAP participants are adjunctively eligible for WIC and they do not need further income verification process. This mechanism increases the choice utility of SNAP participants. Second, SNAP participants are already collecting substantial amount of benefit from SNAP, the additional benefit from WIC has less marginal utility to them as compare to another households that has the same observable characteristics but do not participate in SNAP. This decreases the choice utility of the SNAP-participating households. These two mechanisms likely cancel each other in the choice stage.
Specification (3) adds the racial group indicator for Black and Hispanic households. Black households are less attentive and have a lower choice tendency; Hispanic households are more attentive and have a higher choice tendency. This aligns with the raw data pattern. Black households have a lower than sample average take-up rate, and Hispanic households have a higher than sample average take-up rate. Such racial disparity is persistent even to date, suggesting that outreach to the Black community might be strategic for promoting WIC take-up.
Specification (4) introduces household income in the form of log income per household member as an additional covariate. Counterintuitively, we find that higher income is associated with both greater attention and a higher choice tendency. This aligns with the raw data in Table˜1, which shows that participating households have a higher average income than non-participating households. We hypothesize that this result is driven by car ownership, which likely correlates with income. Vehicle access significantly reduces the hassle costs of traveling to appointments and WIC vendors, potentially outweighing the lower marginal utility of benefits for higher-income households. At the same time, not having a vehicle may make the sign-up process prohibitively troublesome, making households without a car less attentive. While we lack direct data on vehicle ownership to test this mechanism, we find supportive indirect evidence: controlling for income doubles the coefficient on the $S_{it}\times LA_{it}$ interaction. If income proxies for car ownership, this suggests that local accessibility—such as the density of medical services, social assistance offices, and public transit—matters most for households without cars, precisely the group for whom these local amenities are critical for reducing hassle costs.
Specifications (5) and (6) include the full set of covariates: all covariates in the baseline model, SNAP participation status, race, and income. All results remain robust. We interpret the numerical results for specification (6) choice stage in dollar amounts. A household with an infant values the WIC package \$182 more than another without an infant, ceteris paribus. The hassle cost for a household living in a county with no local accessibility is \$224, but each unit increase in $LA_{it}$ reduces hassle cost by more than \$8. Consider a typical household with $LA_{it}=15$, its hassle cost is about \$97. Note that the benefit data is skewed; hence, these dollar amount results are not necessarily interpretable. Next, we consider the attention and choice probabilities under specification (5).
5.2 Heterogeneity in attention and choice probabilities
We first define a baseline household to serve as a benchmark for all our heterogeneity analysis. A baseline household is defined as one infant, no preschooler, in a county where the local accessibility is 15, high school education, non-Black, non-Hispanic, does not participate in SNAP, with an annual income of \$6600 per person in the year 2009. Our model involves a random effect. Hence, we consider a range of random effects $Q_{i}$ ranging from -2 to 2, which covers more than 95% of the standard normal distribution. We will show heterogeneity in both attention and choice probabilities along two dimensions: first, between baseline households and others, second, across different values of $Q$.
Baseline vs no infant
The baseline household has a monthly benefit of \$150, and a household without an infant, all else equal, has a monthly benefit of \$100. We plot the attention probabilities and choice probabilities of the baseline household and a household without an infant but has a preschooler. The rest of the characteristics of these two households are identical. Having an infant is a large attention trigger and increases choice probability substantially, as shown by Figure˜5. Without the infant, households are much less attentive among the lower random effect subpopulation and have a lower choice tendency among the higher random effect subpopulation.
Figure 5: Left panel shows the attention probability of the baseline household and a household without an infant across different values of $Q$; right panel shows the choice probability of the baseline household and a household without an infant across different values of $Q$.
Another dimension of the heterogeneity is across different levels of $Q$. When Q<1Q<-1, the households are hardly attentive, leading to close to huge nonparticipation among this subpopulation. In contrast, as $Q$ increases above 1-1, households become much more attentive. A similar pattern is observed for the choice probability. When $Q>1$, then the choice probability is close to 100%; this subpopulation always chooses participation over nonparticipation. When $Q$ decreases below $1$, then choice probability quickly drops, making these households likely to exit WIC prematurely.
Baseline vs SNAP participants
As shown by the left panel of Figure˜6, the attention probability of a SNAP-participating household is substantially higher than that of the baseline household when Q>1Q>-1. When $0<Q<1$, the difference between the two types of households is as large as 10%. In contrast, the SNAP-participating household’s choice probability is almost identical to the baseline household’s, as shown by the overlapping solid and dashed lines in the right panel of Figure˜6. These two results show that the difference between the take-up rates between SNAP-participating households and other households is mainly driven by attention, not choice.
Figure 6: Left panel shows the attention probability of the baseline household and a household in SNAP across different values of $Q$; right panel shows the choice probability of the baseline household and a household in SNAP across different values of $Q$.
Racial disparity
Figure˜7 displays racial disparities in both attention and choice probabilities. Hispanic households are more attentive and have a higher choice tendency than average. Black households are less attentive and have a lower choice tendency than average. There is more racial disparity in attention probability among households with infants than among households with no infants. The disparity kicks in around Q=1Q=-1 for the top left panel, whereas it kicks in around $Q=0$ for the top right panel. Conversely, there is more racial disparity in choice probability among households with no infants than among households with infants.
Figure 7: Top left panel shows the attention probability of the household by races across different values of $Q$; top right panel shows the attention probability of the household without an infant by races across different values of $Q$. Bottom left panel shows the choice probability of the household by races across different values of $Q$; bottom right panel shows the choice probability of the household without an infant by races across different values of $Q$.
Hassle cost and $LA_{it}$
We compare the choice probability of a household who do not need to pay a hassle, those who need to pay a hassle and stay in a county with $LA_{it}=5$ (LA05), and those who need to pay a hassle and stay in a county with $LA_{it}=15$ (LA15). The comparison shows that hassle cost significantly lowers choice probability when $Q$ is negative. On the other hand, $LA_{it}$ significantly reduces the influence of hassle cost.
Figure 8: Left panel shows the choice probability of the baseline household and a household without an infant across different values of $Q$ when they need to pay a hassle cost to participate in WIC; right panel shows the choice probability of the baseline household and a household without an infant across different values of $Q$ when they need to pay a hassle cost to participate in WIC.
@@ 5.3 Counterfactual simulations
Since our model involves a random effect, we have to simulate households. We divide education into 3 levels: less than high school, high school, and more than high school, denoted as $\{1,2,3\}$. We set $LA_{it}$ as one of the three values: 10, 15, and 20. All households start with one child at the age of $\{3,6,9,12,15\}$ months. Half of the households will have a newborn at the 12th month. In total, there are 90 types of households. For each type, we simulate 5 of them. In total, we simulate 450 households. Each household has a time-invariant random effect drawn from a standard normal. We simulate each household’s decisions for 30 months. We simulate the counterfactuals using specification (1). The general conclusions are robust using different specifications.
@@
Forcing attention increases the participation rate from 47% to 89%. This 42% increase dwarfs the increase from the forcing choice, which is merely 8% (from 47% to 55%). The comparison suggests that attention is the major bottleneck for promoting welfare take-up.
@@
In Figure˜10, we plot the histograms of attention probabilities and choice probabilities of all households under no intervention. The attention probabilities cluster close to 0. Around 48% of the households have their attention probabilities lower than 5%, close to 70% of the households have their attention probabilities lower than 25%. Raising attention probability to 100% induces a huge shift in attention probability’s distribution, so it is extremely effective. In contrast, many households’ choice probabilities cluster close to 1. Around 48% of the households have their attention probabilities higher than 95%, and close to 80% of the households have their attention probabilities higher than 75%. Hence, raising choice probability to 100% is much less effective since many households’ initial choice probabilities are already close to 100%.
@@
This section simulates two realistic policies under the constraint that WIC policymakers can only intervene with currently participating households, as contact information for eligible non-participants is unavailable. Consequently, the only viable strategy is to extend participation spells among enrolled families. The first policy is an attention boost, which raises the attention probability to 100% in the period after a household exits the program. This mimics reminder messages sent to households that exit the program prematurely. The second policy is a choice nudge, which increases the choice probability by a small margin for all participating households with a child under 30 months ($YA_{it}\leq 30$). We test a range of increases, mirroring messages that inform households that continued WIC enrollment is beneficial for the future of the children. Results show that the attention boost raises the participation rate from 46.9% to only 47.4%—a gain of just half a percentage point. In contrast, a choice nudge that increases choice probability by a mere 0.45% achieves the same effect as the attention boost, as shown in Figure˜11.
@@
The effectiveness of choice nudge comes from two mechanisms. One, it directly increases the household’s current period choice probability. Hence, current period retention increases among households who are already enrolled last period. Two, once the households are retained in WIC, they are fully attentive next period due to the ongoing engagement that they have with WIC. This indirectly increases their future attention probability. The combination of the two mechanisms drives the effectiveness of a choice-nudging policy.
Limitations of the Counterfactual Analysis
Our counterfactual analysis indicates that choice-nudging policies may be more effective than attention-boosting interventions, as even a small (0.45%) increase in choice probability can match the effect of a large attention boost. However, this finding rests on a key empirical question: do real-world policies exist that can reliably increase choice probability by such a margin? This limitation—the gap between theoretical policy predictions and their practical implementation—is a common challenge in models of inattention (hortaccsu2017power; heiss2021inattention; einav2025selling). To validate our theoretical prediction, it is necessary to compare the actual treatment effects of attention-boosting and choice-nudging policies in a real-world setting. In Section˜6, we use experimental data from the WIC2Five pilot program to provide exactly this empirical test.
@@ 5.4 Additional findings
On top of showing that choice nudge is likely more effective than attention boost, the full model generates two other interesting empirical findings. First, we find that the salient targeting pattern of choice nudge is on the unobserved random effect, but not on the observables. This is different from the findings from finkelstein2019take, which find a targeting pattern on the observed household characteristics. We provide a unifying explanation for these two targeting patterns. Second, we note that higher education is associated with lower attention but higher choice tendency. This finding opens a scope for heterogeneous policy targeting. We defer the details of these additional findings to Appendix˜E.